The null hypothesis regarding a contingency table with 6 rows and 4 columns should be rejected (at 95% confidence) if the test statistic is greater than


1
Expert's answer
2022-01-11T13:50:25-0500

Degrees of freedom: "df=(c-1)(r-1)=(4-1)(6-1)=15."

For "\alpha=0.05, \;df=15:\chi^2_{crit}=24.996."

The null hypothesis regarding a contingency table should be rejected (at 95% confidence) if the test statistic is greater than 24.996.